# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 2. Preliminary results
$(a+b)^{\wp}\leq(1+\eta)a^{\wp}+C_{\eta}b^{\wp},\quad\forall~{}a,b>0,$ (3.6)
where $C_{\eta}=(1-(1+\eta)^{-\frac{1}{\wp-1}})^{1-\wp}$. Moreover, by using (3.6) and Young’s inequality, we have
$\int_{S_{2}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$ $\leq\int_{S_{2}}\frac{\operatorname{\exp}(\alpha|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{S_{2}\cap~{}\textit{supp}(\psi)}\frac{\operatorname{\exp}(\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}}+C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{S_{2}\cap~{}\textit{supp}(\psi)}\frac{\operatorname{\exp}(2\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\textit{supp}(\psi)}\frac{\operatorname{\exp}(2C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x.$
Let $\eta,\varepsilon>0$ be sufficiently small enough such that $2\alpha(1+\eta)\varepsilon^{\frac{n}{n-2}}<\beta_{\gamma,n}$ holds. Hence we obtain by using the definition of $S_{2}$ and Theorem 1.3 that
$\int_{S_{2}\cap~{}\textit{supp}(\psi)}\frac{\operatorname{\exp}(2\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$ $\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{2\alpha(1+\eta),j_{0}}(u-\psi)}{|x|^{\gamma}}\,\mathrm{d}x$
@@
In addition, since $\psi$ has compact support, we have
$\int_{\textit{supp}(\psi)}\frac{\operatorname{\exp}(2C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
Thus it follows that
$\int_{S_{2}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$ (3.7)
In view of (3.5) and (3.7), we conclude that the statement of the corollary is well established. ∎
The proof of Theorem 1.4 is based on the following crucial lemma. More precisely, we have
@@
$\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,\quad\forall~{}\ell^{\prime}\in(\ell,L_{n}(u)).$
@@
By the standard density argument, we can assume a sequence $\{\psi_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u_{k}-\psi_{k}\|<\frac{1}{k},~{}\forall~{}k\in\mathbb{N}$. Suppose that $\overset{\sim}{u}_{k}:=\frac{\psi_{k}}{\|\psi_{k}\|},~{}\forall~{}k\in\mathbb{N}$. It follows that the sequence $\{\overset{\sim}{u}_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ and satisfying
$\|\overset{\sim}{u}_{k}\|=1,\quad\|u_{k}-\overset{\sim}{u}_{k}\|\leq\frac{2}{k},~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad\overset{\sim}{u}_{k}\rightharpoonup u\quad\text{in}\quad E\quad\text{as}\quad k\to\infty.$ (3.8)
@@
$J_{1}$ $=\int_{S_{1}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})^{j}}{j!}\Big{|}\frac{u_{k}}{2}\Big{|}^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{S_{1}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})^{j}}{j!}\Big{|}\frac{u_{k}}{2}\Big{|}^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|<1\}}\frac{|u_{k}|^{p^{\ast}}}{2^{p^{\ast}}|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|\geq 1\}}\frac{|u_{k}|^{p^{\ast}}}{2^{p^{\ast}}|x|^{\gamma}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\int_{\{x\in S_{1}\colon~{}|x|\geq 1\}}|u_{k}|^{p^{\ast}}\,\mathrm{d}x\leq C(\gamma,\ell,n),$ (3.9)
where $C(\gamma,\ell,n)>0$ is a constant depending only on $\gamma,\ell$ and $n$ but independent on $k$. Further, we can also write
@@
By direct calculation, one has $|\overset{\sim}{u}_{k}(x)|>1$ for all $x\in\Gamma_{1}$ and $k\in\mathbb{N}$. In virtue of (3.6) with $\eta=\frac{\ell^{\prime}-\ell}{\ell}$, we infer that
$J_{2,\Gamma_{1}}$ $\leq\int_{\Gamma_{1}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{1}}\frac{\operatorname{\exp}(\ell^{\prime}\beta_{\gamma,n}|\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{\eta}\ell\beta_{\gamma,n})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$ (3.10)
@@
Utilizing the inequality (3.6) with $\eta=1$, we obtain from the definition of $\Gamma_{2}$ that
$J_{2,\Gamma_{21}}$ $\leq\int_{\Gamma_{21}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{21}}\frac{\operatorname{\exp}(2\ell\beta_{\gamma,n}|u_{k}-\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{1}\ell\beta_{\gamma,n})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$ (3.11)
@@
$\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C,~{}\forall~{}k\in\mathbb{N}.$ (3.12)
From (3) and (3.12), one can notice that
$J_{2,\Gamma_{21}}=\int_{\Gamma_{21}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C,~{}\forall~{}k\in\mathbb{N},$ (3.13)
@@
$J_{2,\Gamma_{22}}$ $\leq\int_{\Gamma_{22}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{22}}\frac{\operatorname{\exp}((1+\eta)\ell\beta_{\gamma,n}|\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{\eta}\ell\beta_{\gamma,n}|u_{k}-\overset{\sim}{u}_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
@@
$\leq C_{1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+C_{2}\int_{\mathbb{R}^{n}}\frac{\Phi_{t^{\prime}C_{\eta}\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,$ (3.14)
where $C_{1},~{}C_{2}>0$ are some suitable constants independent on the choice of $k$. Let $k_{0}\in\mathbb{N}$ be sufficiently large enough such that $t^{\prime}C_{\eta}\ell\beta_{\gamma,n}\big{(}\frac{2}{k}\big{)}^{\frac{n}{n-2}}<\beta_{\gamma,n}$ for all $k\geq k_{0}$. Then arguing as before, it is not difficult to prove that
$\int_{\mathbb{R}^{n}}\frac{\Phi_{t^{\prime}C_{\eta}\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C_{3},~{}\forall~{}k\in\mathbb{N},$ (3.15)
for some suitable constant $C_{3}>0$ independent of $k$. Thus by using (3) and (3.15), we infer that
$\sup_{k\in\mathbb{N}}J_{2,\Gamma_{22}}\leq C_{1}\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+C_{4},$ (3.16)
where $C_{4}>0$ is some suitable constant independent on the choice of $k$. Hence by using (3), (3), (3.13) and (3.16), we conclude the proof. ∎
@@
The proof of the first part of Theorem 1.4 is based on method of contradiction. Due to Lemma 3.2, it suffices to prove Theorem 1.4 for compactly supported smooth functions. Indeed, let $\{u_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ be such that for any $\ell\in(0,L_{n}(u))$, we have
$\|u_{k}\|=1,\quad u_{k}\rightharpoonup u\neq 0\quad\text{in}\quad E\quad\text{as}\quad k\to\infty\quad\text{and}\quad\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\ \mathrm{d}x=+\infty.$ (3.17)
Let $u_{k}^{\ast}$ be the Schwarz symmetrization of $u_{k}$. Notice that $(1/|x|^{\gamma})^{\ast}=1/|x|^{\gamma}$ and therefore by applying Hardy-Littlewood inequality, for instance, see Lemma 2.1 and (2.2), we have
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\ \mathrm{d}x\leq\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$ (3.18)
@@
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\nearrow\infty\quad\text{as}\quad k\to\infty.$ (3.19)
@@
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x=\int_{B_{R}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{B_{R}^{c}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,\quad\forall~{}k\in\mathbb{N}.$
@@
where $c_{1}>0$ is a suitable constant independent on the choice of $k$. Now by using Lemma 2.4, Lemma 2.5 and (3.17), one sees that
$|u_{k}^{\ast}(x)|\leq\frac{1}{|x|^{\frac{n}{p^{\ast}}}}\Big{(}\frac{n}{\omega_{n-1}}\Big{)}^{\frac{1}{p^{\ast}}}\|u_{k}\|_{p^{\ast}}\leq\Big{(}\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}}\Big{)}^{\frac{1}{p^{\ast}}}\frac{1}{|x|^{\frac{n}{p^{\ast}}}},~{}\forall~{}~{}x\neq 0,$
@@
$\int_{B_{R}^{c}}|u^{\ast}_{k}|^{\frac{nj}{n-2}}\,\mathrm{d}x$ $\leq\omega_{n-1}\Big{(}\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}}\Big{)}^{\frac{nj}{(n-2)p^{\ast}}}\int_{R}^{\infty}r^{n-\frac{n}{p^{\ast}}\frac{nj}{n-2}-1}\,\mathrm{d}r$
@@
On combining all the above information, one has that
@@
$:=C(\ell,\gamma,R,n)<+\infty,$
where $C(\ell,\gamma,R,n)>0$ is a constant depending only on $\ell,\gamma,R$ and $n$. In conclusion, we infer that
$\sup_{k\in\mathbb{N}}\int_{B_{R}^{c}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C(\ell,\gamma,R,n)<+\infty.$ (3.20)

### v3

@@ 2. Preliminary results
(3.6) $(a+b)^{\wp}\leq(1+\eta)a^{\wp}+C_{\eta}b^{\wp},~\forall~a,b>0,$
where $C_{\eta}=\big(1-(1+\eta)^{-\frac{1}{\wp-1}}\big)^{1-\wp}$. Moreover, by using (3.6) and Young’s inequality, one has
$\int_{S_{2}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x$ $\leq\int_{S_{2}}\frac{\operatorname{\exp}(\alpha|u|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{S_{2}\cap~\text{supp}(\psi)}\frac{\operatorname{\exp}(\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}}+C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{S_{2}\cap~\text{supp}(\psi)}\frac{\operatorname{\exp}(2\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\text{supp}(\psi)}\frac{\operatorname{\exp}(2C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x.$
Let $\eta,\varepsilon>0$ be sufficiently small enough such that $2\alpha(1+\eta)\varepsilon^{\frac{n}{n-2}}<\beta_{\gamma,n}$ holds. Hence, we obtain by using the definition of $S_{2}$ and Theorem 1.1 that
$\int_{S_{2}\cap~\text{supp}(\psi)}\frac{\operatorname{\exp}(2\alpha(1+\eta)|u-\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$ $\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{2\alpha(1+\eta),j_{0}}(u-\psi)}{|x|^{\gamma}}\,\mathrm{d}x$
@@
By using $\psi$ has compact support, we obtain
$\int_{\text{supp}(\psi)}\frac{\operatorname{\exp}(2C_{\eta}\alpha|\psi|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
This immediately implies that
(3.7) $\int_{S_{2}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
Based on (3.5) and (3.7), we infer that the assertion of the corollary is well-established. ∎
4. Sharp singular concentration-compactness principle: Proof of Theorem 1.2
This section focuses on proving the concentration-compactness principle of the singular Adams’ type inequality stated in the previous section (see Theorem 1.1), which relies on conventional Schwarz symmetrization techniques. More information for a detailed analysis in this direction can be found in [23, 44]. The proof of Theorem 1.2 is based on the following crucial lemma. More precisely, we have
@@
$\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,~\forall~\ell^{\prime}\in(\ell,L_{n}(u)).$
@@
By the standard density argument, we can assume a sequence $\{\psi_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u_{k}-\psi_{k}\|<\frac{1}{k}$ for all $k\in\mathbb{N}$. Suppose that $\overset{\sim}{u}_{k}:=\frac{\psi_{k}}{\|\psi_{k}\|}$ for all $k\in\mathbb{N}$. It follows that the sequence $\{\overset{\sim}{u}_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ and satisfying
(4.1) $\|\overset{\sim}{u}_{k}\|=1,\quad\|u_{k}-\overset{\sim}{u}_{k}\|\leq\frac{2}{k},~\forall~k\in\mathbb{N}\quad\text{and}\quad\overset{\sim}{u}_{k}\rightharpoonup u\quad\text{in}\quad E\quad\text{as}\quad k\to\infty.$
@@
(4.2) $J_{1}$ $=\int_{S_{1}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})^{j}}{j!}\Big|\frac{u_{k}}{2}\Big|^{\frac{nj}{n-2}}\,\mathrm{d}x\leq\int_{S_{1}}\frac{1}{|x|^{\gamma}}\sum_{j=j_{0}}^{\infty}\frac{(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})^{j}}{j!}\Big|\frac{u_{k}}{2}\Big|^{p^{\ast}}\,\mathrm{d}x$
$\leq\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\Bigg[\int_{\{x\in S_{1}\colon~|x|<1\}}\frac{|u_{k}|^{p^{\ast}}}{2^{p^{\ast}}|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in S_{1}\colon~|x|\geq 1\}}\frac{|u_{k}|^{p^{\ast}}}{2^{p^{\ast}}|x|^{\gamma}}\,\mathrm{d}x\Bigg]$
$\leq\operatorname{\exp}(2^{\frac{n}{n-2}}\ell\beta_{\gamma,n})\Bigg[\int_{\{x\in S_{1}\colon~|x|<1\}}\frac{1}{|x|^{\gamma}}\,\mathrm{d}x+\int_{\{x\in S_{1}\colon~|x|\geq 1\}}|u_{k}|^{p^{\ast}}\,\mathrm{d}x\Bigg]$
$\leq C(\gamma,\ell,p,n),$
where $C(\gamma,\ell,p,n)>0$ is a constant depending only on $\gamma,\ell,p$ and $n$ but independent on $k$. Further, we can also write
@@
By direct calculation, one has $|\overset{\sim}{u}_{k}(x)|>1$ for all $x\in\Gamma_{1}$ and $k\in\mathbb{N}$. In virtue of (3.6) with $\eta=\frac{\ell^{\prime}-\ell}{\ell}$, we get
(4.3) $J_{2,\Gamma_{1}}$ $\leq\int_{\Gamma_{1}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{1}}\frac{\operatorname{\exp}(\ell^{\prime}\beta_{\gamma,n}|\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{\eta}\ell\beta_{\gamma,n})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\lesssim\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$
@@
By utilizing the inequality (3.6) with $\eta=1$, we obtain from the definition of $\Gamma_{2}$ that
(4.4) $J_{2,\Gamma_{21}}$ $\leq\int_{\Gamma_{21}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{21}}\frac{\operatorname{\exp}(2\ell\beta_{\gamma,n}|u_{k}-\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{1}\ell\beta_{\gamma,n})}{|x|^{\gamma}}\,\mathrm{d}x$
$\lesssim\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$
@@
(4.5) $\int_{\mathbb{R}^{n}}\frac{\Phi_{2\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C,~\forall~k\in\mathbb{N}.$
From (4.4) and (4.5), one can notice that
(4.6) $J_{2,\Gamma_{21}}=\int_{\Gamma_{21}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C,~\forall~k\in\mathbb{N},$
@@
(4.7) $J_{2,\Gamma_{22}}$ $\leq\int_{\Gamma_{22}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x\leq\int_{\Gamma_{22}}\frac{\operatorname{\exp}((1+\eta)\ell\beta_{\gamma,n}|\overset{\sim}{u}_{k}|^{\frac{n}{n-2}}+C_{\eta}\ell\beta_{\gamma,n}|u_{k}-\overset{\sim}{u}_{k}|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x$
@@
$\leq C_{1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+C_{2}\int_{\mathbb{R}^{n}}\frac{\Phi_{t^{\prime}C_{\eta}\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,$
where $C_{1}>0$ and $C_{2}>0$ are some suitable constants independent on the choice of $k$. Let $k_{0}\in\mathbb{N}$ be sufficiently large enough such that $t^{\prime}C_{\eta}\ell\beta_{\gamma,n}\big(\frac{2}{k}\big)^{\frac{n}{n-2}}<\beta_{\gamma,n}$ for all $k\geq k_{0}$. Then arguing as before, it is not difficult to prove that
(4.8) $\int_{\mathbb{R}^{n}}\frac{\Phi_{t^{\prime}C_{\eta}\ell\beta_{\gamma,n},j_{0}}(u_{k}-\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C_{3},~\forall~k\in\mathbb{N},$
for some suitable constant $C_{3}>0$ independent of $k$. Thus, by using (4.7) and (4.8), we infer that
(4.9) $\sup_{k\in\mathbb{N}}J_{2,\Gamma_{22}}\leq C_{1}\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell^{\prime}\beta_{\gamma,n},j_{0}}(\overset{\sim}{u}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+C_{4},$
where $C_{4}>0$ is some suitable constant independent of the choice of $k$. Hence, by using (4.2), (4.3), (4.6) and (4.9), we conclude the proof. ∎
Now, we are in a position to prove the concentration-compactness principle as follows.
@@
To prove Theorem 1.2, first one can see that $\Phi_{\alpha,j_{0}}(|u|)=\Phi_{\alpha,j_{0}}(u)$ and thus, without loss of generality, we assume that $\{u_{k}\}_{k}\subset E$ and its weak limit $u$ are nonnegative functions. Moreover, it follows from the weak lower semicontinuity of the norm in $E$ that
$\|u\|\leq\liminf_{k\to\infty}\|u_{k}\|=1.$
Thus, the proof of Theorem 1.2 is based on two cases as follows:
Case 1: In this case, we consider the situation $\|u\|=1$. Recall that $E$ is a reflexive Banach space, $u_{k}\rightharpoonup u$ in $E$ and $\|u_{k}\|\to\|u\|$ in $\mathbb{R}$ as $k\to\infty$. Hence, one can deduce from [6, Corollary A.2] that $u_{k}\to u$ in $E$ as $k\to\infty$. Further, by using similar arguments used in [55, Lemma 3.2], up to a subsequence of $\{u_{k}\}_{k}\subset E$ still denoted by same symbol, there exists some $w\in E$ such that $|u_{k}(x)|\leq w(x)$ for a.e. $\mathbb{R}^{n}$ for all $k\in\mathbb{N}$. Thus, we obtain from Corollary 3.1 and the Lebesgue dominated convergence theorem that
(4.10) $\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x=\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
Moreover, by using Corollary 3.1 and (4.10), we conclude that (1.12) holds.
Case 2: In this case, we consider the situation $0<\|u\|<1$. The proof of the first part of Theorem 1.2 is based on the method of contradiction. Due to Lemma 4.1, it suffices to prove Theorem 1.2 for compactly supported smooth functions. Indeed, let $\{u_{k}\}_{k}\subset C^{\infty}_{0}(\mathbb{R}^{n})$ be such that for any $\ell\in(0,L_{n}(u))$, we have
(4.11) $\|u_{k}\|=1,\quad u_{k}\rightharpoonup u\neq 0\quad\text{in}\quad E\quad\text{as}\quad k\to\infty\quad\text{and}\quad\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\ \mathrm{d}x=+\infty.$
Let $u_{k}^{\ast}$ be the Schwarz symmetrization of $u_{k}$. Notice that $(1/|x|^{\gamma})^{\ast}=1/|x|^{\gamma}$ and therefore, by applying Hardy-Littlewood inequality, for instance, see Lemma 2.1 and (2.2), we have
(4.12) $\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\ \mathrm{d}x\leq\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x.$
@@
(4.13) $\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\nearrow+\infty\quad\text{as}\quad k\to\infty.$
@@
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x=\int_{B_{R}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x+\int_{B_{R}^{c}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x,~\forall~k\in\mathbb{N}.$
@@
where $c_{1}>0$ is a suitable constant independent of the choice of $k$. Now, by using Lemma 2.4, Lemma 2.5 and (4.11), one sees that
$|u_{k}^{\ast}(x)|\leq\frac{1}{|x|^{\frac{n}{p^{\ast}}}}\bigg(\frac{n}{\omega_{n-1}}\bigg)^{\frac{1}{p^{\ast}}}\|u_{k}\|_{p^{\ast}}\leq\bigg(\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}}\bigg)^{\frac{1}{p^{\ast}}}\frac{1}{|x|^{\frac{n}{p^{\ast}}}},~\forall~~x\neq 0,$
@@
$\int_{B_{R}^{c}}|u^{\ast}_{k}|^{\frac{nj}{n-2}}\,\mathrm{d}x$ $\leq\omega_{n-1}\bigg(\frac{nc_{2}^{p^{\ast}}}{\omega_{n-1}}\bigg)^{\frac{nj}{(n-2)p^{\ast}}}\int_{R}^{\infty}r^{n-\frac{n}{p^{\ast}}\frac{nj}{n-2}-1}\,\mathrm{d}r$
@@
By combining all the above information, one has
@@
$:=C(\ell,\gamma,R,p,n)<+\infty,$
where $C(\ell,\gamma,R,p,n)>0$ is a constant depending only on $\ell,\gamma,R,p$ and $n$. In conclusion, we infer that
(4.14) $\sup_{k\in\mathbb{N}}\int_{B_{R}^{c}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u^{\ast}_{k})}{|x|^{\gamma}}\,\mathrm{d}x\leq C(\ell,\gamma,R,p,n)<+\infty.$